Problem Solving Flowchart

Anonymous September 3, 1752

Begin Here Does It Work?

Don’t Mess With It!

Did You Mess With It?

You Idiot! Does Anyone Know?

Will You Pay For It?

Will You Be Fired?

You Poor Fool! Can You Blame Someone Else?

NO PROBLEM!

Hide It!

Yes

No

Yes

No

Yes

Yes Yes

No

No

No Yes No

Figure 1: Problem Solving Flowchart